A 70.0 kg weight sits at the top of a ramp that makes an angle of 40o with the horizontal. As the weight slides down the ramp, it experiences a force of friction of 75 N. What is the acceleration of the weight down the ramp?


1
Expert's answer
2021-02-08T18:42:29-0500

From the Newton's second law:


"ma=mgsin\alpha-F_{fr},"


where m - the mass of a weight;

"\alpha" - angle, that the ramp makes with the horizontal;

"F_{fr} -" a force of friction.


Hence,


"a=\frac{mgsin\alpha-F_{fr}}{m}=\frac{70\cdot 9.81\cdot sin40\degree-75}{70}=5.23\space m/s^2."


Answer: 5.23 "m\/s^2."